OpenStudy (anonymous):

Factor 4m^16 - 12m^8p^3 + 9p^6 completely.

OpenStudy (yttrium):

Do you remember this algebraic form? \[(a+b)^2 = a^2 + 2ab + b^2\]

OpenStudy (yttrium):

Well, another form is; \[(a-b)^2 = a^2 - 2ab +b^2\]